On the Apollo 14 mission to the moon, astronaut Alan Shepard hit a golf ball with a 6 iron. The acceleration due to gravity on the moon is 1/6 of its value on earth. Suppose he hits the ball with a speed of 14 m/s at an angle 55 ∘ above the horizontal.

Part A

How much farther did the ball travel on the moon than it would have on earth?

Express your answer using two significant figures.

Part B

For how much more time was the ball in flight on the moon than it would have on earth?

Express your answer using two significant figures.
Answer #1

Given that velocity u=14 m/s

angle θ=55 deg

accleration due to gravity on the moon is gm= 1/6(g)

Find flight time on Earth t(earth)

= 2vi sin theta / g(earth)

=2*14m/s*sin55° / 9.8 m/s^2

=2.34 s

Find flight time on moon

time of flight t=2usin tetha /gm

                          =2 * 14 * sin 55 / (9.8/6)

=22.936 / 1.6333

                          =14.04 s

the distance travelled on moon is

Rm=u^2 Sin 2theta /gm

=14^2 sin(2*55) / (9.8/6)

      =112.76 m

range on earth is

R=u^2 sin2 theta /g

=14^2 sin (2*55) / 9.8

      =18.79 m

so extra distance travalled is S=Rm-R

=112.76-18.79=93.97 m

I hope help you !!